Bank One has loans worth $350, reserves of $60, checkable deposits at $390 and capital of $20. If the loan performance deteriorates to a bad economy, what is the maximum write down in loan value allowed, such that the bank's ratio of capital to total assets does not fall below 2%?

$9.5

$12

$13.3

$14.4
Answer #1

Total assets = 350+60 = 410
Let loan amount write down = x
Bank's Ratio of capital to Total assets after write down = (20-x)/(410-x) = 2%
20 -x = 410*2% -2%x
20-8.20 = 98x
x =11.80/0.98 = 12.04 or 12 (Option b is correct option)

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well